The principal P is borrowed at a simple interest rate r for a period of time t. Find the simple interest owed for the use of the money. Assume 360 days in a year.
P​ = ​$​230, r​ = ​3%, t​ = 2 years

Answer by ikleyn(52781)

.

Simple interest  I = P*r*t = 230*0.03*2  dollars.


The number of days (360) is not relevant to the solution of the given problem.

---------------

On simple interest problems,  see the lesson
    - Simple interest percentage problems
in this site.